THIS IS THE MOST IMPORTANT QUESTION ON MY HOMEWORK AND I JUST CAN'T FIGURE IT OUT!!! please help !
A 25 foot ladder is placed against a building. The bottom of the ladder is 7 feet from the building. If the top of the ladder slips down by 4 feet, by how many feet will the bottom slide out?

Answer by scott8148(6628)

let x=distance bottom slides out, let y=height of top

by Pythagoras __ 7^2+y^2=25^2 __ y^2=576 __ y=24

"the top of the ladder slips down by 4 feet" __ y-4=20

(7+x)^2+20^2=25^2 __ (7+x)^2=225 __ 7+x=15 __ x=8